# NeurOp-Diff: Continuous Remote Sensing Image Super-Resolution via Neural Operator Diffusion

### v2

@@ 3.4 Network details
NeurOp-Diff network architecture is illustrated in the figure 2. The neural operator utilizes a scaling factor $s\sim\mathcal{U}(1,M]$ to ensure the model is effectively applicable to multiple scales. In this experiment, we set $M=8$. For the feature encoder $E$, we reference the EDSR-baseline [29] by removing the last sampling layer. Regarding the noise level $\gamma$, we follow a uniform sampling method [34]. First, the time step $t\sim\{0,T\}$ is sampled uniformly, followed by sampling $\gamma$ from $\mathcal{U}(\gamma_{t-1},\gamma_{t})$. This process is expressed as: $\gamma\sim\frac{1}{T}\sum_{t=1}^{T}\mathcal{U}(\gamma_{t-1},\gamma_{t})$.
@@ 4.1 Training
In this experiment, we used three remote sensing datasets: UCMerced [46], AID [44] and RSSCN7 [52]. A portion of each dataset was used as the training set, while the remaining portion was used for evaluation. For AID and RSSCN7, which do not match the target resolution, we adjusted the data using center cropping [3] and downsampling. We compared NeurOp-Diff with various representative super-resolution methods, including both diffusion-based generative approaches and regression-based continuous super-resolution methods. Among the generative methods, SR3 [34] was used as a baseline and reference, and comparisons were made with IDM [11], EDiffSR [45], and TCDM [47]. For regression-based methods, LIIF [6] was used as a reference, and comparisons were made with SRNO [42]. To evaluate the effectiveness of our continuous super-resolution method, we also tested magnification factors beyond the training distribution when compared with SRNO. During the training phase, the scaling factors were uniformly sampled within a continuous range from ×1 to ×8. In the testing phase, the model was further evaluated on images with out-of-distribution higher scaling factors, specifically from ×9 to ×10. The experimental results are primarily evaluated using PSNR (Peak Signal-to-Noise Ratio) and SSIM (Structural Similarity Index).
@@
Following the referenced approach [42], we first pre-trained the neural operator component for 500 epochs with a batch size of 64. Then, we froze the weights of the neural operator and trained the entire NeurOp-Diff model. The batch size was set to 10, and approximately 1 M iterations were performed. The Adam optimizer [23] was used, with an initial learning rate of $1\times 10^{-4}$ and a minimum learning rate of $2\times 10^{-6}$. For the first 0.1 M iterations, a constant learning rate was maintained, followed by a learning rate decay strategy. Dropout was set to 0.2 during training.
During training, the scaling factor $s\sim\mathcal{U}(1,M]$ was sampled from a uniform distribution to cover multiple levels of magnification. For super-resolution tasks, a larger $T$ generally yields better results. Therefore, in this experiment, we set $T=2000$. And we chose the $L_{1}$ norm as the loss function. In the inference phase, we used a uniform sampling strategy [36] to accelerate inference. Experiments showed that setting the number of noise diffusion steps to 50 produced good inference results. The experiments were conducted on an NVIDIA RTX 4090 GPU.
@@
UCMerced SR3 [34] 27.15 0.7587 23.12 0.5846
IDM [11] 27.66 0.7606 23.68 0.6131
TCDM [47] 27.83 0.7679 24.17 0.6422
@@
AID SR3 [34] 26.24 0.6709 22.54 0.5231
IDM [11] 27.03 0.6778 22.90 0.5487
EDiffSR [45] 27.46 0.6821 23.22 0.5562
@@
RSSCN7 SR3 [34] 26.22 0.5835 22.96 0.5329
IDM [11] 27.18 0.6235 23.37 0.5603
EDiffSR [45] 27.53 0.6389 23.58 0.5827
@@
Table 1: Quantitative comparison of various generative super-resolution ($64\times 64\rightarrow 256\times 256$) methods on UCMerced, AID, and RSSCN7 in terms of average PSNR and SSIM. The best performance are bolded.
@@ 4.2.1 Qualitative Comparison
Figure 3 presents the super-resolution results of $64\times 64\to 256\times 256$ for the UCMerced and AID remote sensing datasets. Compared with various state-of-the-art (SOTA) generative methods, NeurOp-Diff demonstrates excellent reconstruction capability. Although all methods can generally produce realistic super-resolution images, the magnified regions marked with rectangles reveal that NeurOp-Diff demonstrates superior reconstruction performance in recovering high-frequency components such as road textures and shadows, compared to other generative methods. This is attributed to the prior knowledge provided by the neural operator.
@@ 4.2.2 Quantitative Comparison
We evaluated NeurOp-Diff on three remote sensing datasets: UCMerced, AID, and RSSCN. Table 1 presents the PSNR and SSIM results of various models for 4x SR and 8x SR. Compared to other generative models, NeurOp-Diff demonstrates superior performance in both pixel-level accuracy (PSNR) and visual perceptual quality (SSIM).
@@
LIIF [6] 28.96 0.69 0.182
SRNO [42] 29.21 0.72 0.164
@@
Table 3: Quantitative comparison of continuous super-resolution (SR) results on UCMerced (PSNR/SSIM).
Method in-distribution out-of-distribution
2×2$\times$2 × 3.1×3.1$\times$3.1 × 5.8×5.8$\times$5.8 × 7×7$\times$7 × 8×8$\times$8 × 9×9$\times$9 × 10×10$\times$10 ×
SRNO [42] 34.32/0.786 31.05/0.746 27.19/0.681 25.73/0.649 24.67/0.623 24.32/0.607 23.92/0.592
NeurOp-Diff 33.93/0.824 30.76/0.792 26.89/0.717 25.50/0.673 24.51/0.650 24.13/0.641 23.88/0.633
4.3 Comparison of Continuous SR
Table 2 presents the quantitative comparison of LIIF, SRNO, and NeurOp-Diff on UCMerced. Although LIIF and SRNO achieve higher PSNR, our NeurOp-Diff demonstrates better performance in terms of SSIM. Moreover, in subjective perceptual differences (LPIPS), the images generated by our model show smaller differences compared to the real images. Furthermore, as shown in Figure 4, regression-based continuous super-resolution methods struggle to reconstruct high-frequency details and restore realistic textures, often producing overly smooth images. For instance, fine ground stripes are almost invisible in the images output by SRNO. Even when reconstructing more prominent stripes, the results appear somewhat blurry. In contrast, NeurOp-Diff demonstrates superior capability in restoring these details.
From a quantitative perspective (see Table 3), although SRNO demonstrates slightly better PSNR values at lower magnification factors (e.g., 2x to 3.1x), NeurOp-Diff demonstrates superior performance in terms of SSIM metrics, which better aligns with human visual perception. As the magnification factor increases, the advantages of NeurOp-Diff become more apparent, particularly in challenging scenarios with higher magnification factors (e.g., 7x to 10x), where it consistently maintains higher SSIM values. Visually, NeurOp-Diff exhibits better performance in terms of the clarity of tennis court boundary lines and natural texture transitions in the surrounding environment. These advantages become more pronounced at higher magnification factors. This is primarily because NeurOp-Diff is better at capturing and reconstructing high-frequency information, allowing the generated images at high magnification factors to maintain high visual quality and detail richness. However, SRNO often struggles to fully recover complex texture details when handling larger magnifications, resulting in relatively lacking detail in the generated images.
@@ 4.4 Ablation Studies
To explore the impact of the LR conditional network, we conducted 4× super-resolution ablation experiments on remote sensing datasets. Specifically, we constructed comparison models by replacing the scale-adaptive conditional network in NeurOp-Diff with three different conditional mechanisms: (1) Concatenating the upsampled LR image directly with the ground truth image (SR3 [34]); (2) Concatenating LR features encoded by EDSR [29] with the ground truth image; (3) Concatenating LR features encoded by the neural operator with the ground truth image. As shown in Figure 1, directly using the upsampled LR image as a prior condition results in the least amount of detail recovery. Encoding LR features with EDSR recovers more details, but it still falls short compared to the scale-adaptive network encoded by the neural operator. In comparison, it is evident that LR images encoded by the neural operator provide more prior knowledge.

### v3

@@ 3.4 Network details
NeurOp-Diff network architecture is illustrated in the figure 2. The neural operator utilizes a scaling factor $s\sim\mathcal{U}(1,M]$ to ensure the model is effectively applicable to multiple scales. In this experiment, we set $M=8$. For the feature encoder $E$, we reference the EDSR-baseline [Lim2017] by removing the last sampling layer. Regarding the noise level $\gamma$, we follow a uniform sampling method. First, the time step $t\sim\{0,T\}$ is sampled uniformly, followed by sampling $\gamma$ from $\mathcal{U}(\gamma_{t-1},\gamma_{t})$. This process is expressed as: $\gamma\sim\frac{1}{T}\sum_{t=1}^{T}\mathcal{U}(\gamma_{t-1},\gamma_{t})$ [saharia2022image].
@@ 4.1 Training
In this experiment, we used three remote sensing datasets: UCMerced [yang2010bag], AID [xia2017aid] and RSSCN7 [zou2015]. A portion of each dataset was used as the training set, while the remaining portion was used for evaluation. For AID [xia2017aid] and RSSCN7 [zou2015], which do not match the target resolution, we adjusted the data using center cropping [brock2018large] and downsampling. We compared NeurOp-Diff with various representative super-resolution methods, including both diffusion-based generative approaches and regression-based continuous super-resolution methods. To evaluate the effectiveness of our continuous super-resolution method, we also tested magnification factors beyond the training distribution when compared with SRNO [wei2023super]. During the training phase, the scaling factors were uniformly sampled within a continuous range from ×1 to ×8. In the testing phase, the model was further evaluated on images with out-of-distribution higher scaling factors, specifically from ×9 to ×10. The experimental results are primarily evaluated using PSNR (Peak Signal-to-Noise Ratio) and SSIM (Structural Similarity Index) [wang2004image].
@@
Following the referenced approach [wei2023super], we first pre-trained the neural operator component for 500 epochs with a batch size of 64. Then, we froze the weights of the neural operator and trained the entire NeurOp-Diff model. The batch size was set to 10, and approximately 1M iterations were performed. The Adam optimizer [kingma2014adam] was used, with an initial learning rate of $1\times 10^{-4}$ and a minimum learning rate of $2\times 10^{-6}$. For the first 0.1 M iterations, a constant learning rate was maintained, followed by a learning rate decay strategy. Dropout was set to 0.2 during training. During training, the scaling factor $s\sim\mathcal{U}(1,M]$ was sampled from a uniform distribution to cover multiple levels of magnification. For super-resolution tasks, a larger $T$ generally yields better results. Therefore, in this experiment, we set $T=2000$. And we chose the $L_{1}$ norm as the loss function. In the inference phase, we used a uniform sampling strategy [song2020denoising] to accelerate inference. Experiments showed that setting the number of noise diffusion steps to 50 produced good inference results. The experiments were conducted on an NVIDIA RTX 4090 GPU.
@@
UCMerced SR3 [saharia2022image] 27.15 0.7587 23.12 0.5846
IDM [gao2023implicit] 27.66 0.7606 23.68 0.6131
TCDM [zhang2024] 27.83 0.7679 24.17 0.6422
@@
AID SR3 [saharia2022image] 26.24 0.6709 22.54 0.5231
IDM [gao2023implicit] 27.03 0.6778 22.90 0.5487
TCDM [zhang2024] 27.53 0.6840 23.37 0.5584
@@
RSSCN7 SR3 [saharia2022image] 26.22 0.5835 22.96 0.5329
IDM [gao2023implicit] 27.18 0.6235 23.37 0.5603
TCDM [zhang2024] 27.71 0.6489 23.69 0.5873
@@
Table 1: Quantitative comparison of various generative super-resolution ($64\times 64\rightarrow 256\times 256$, $32\times 32\rightarrow 256\times 256$) methods on UCMerced, AID, and RSSCN7 in terms of average PSNR and SSIM. The best performance are bolded.
@@ 4.2.1 Qualitative Comparison
Figure 3 presents the super-resolution results of $64\times 64\to 256\times 256$ for the UCMerced and AID remote sensing datasets. Compared with various state-of-the-art (SOTA) generative methods, NeurOp-Diff demonstrates excellent reconstruction capability. Although all methods can generally produce realistic super-resolution images, the magnified regions marked with rectangles reveal that NeurOp-Diff demonstrates superior reconstruction performance in recovering high-frequency components such as road textures and shadows, compared to other generative methods. This is attributed to the prior knowledge provided by the NO.
@@ 4.2.2 Quantitative Comparison
Table 1 presents the PSNR and SSIM results of various models for 4x SR and 8x SR. Compared to other generative models, NeurOp-Diff demonstrates superior performance in both pixel-level accuracy (PSNR) and visual perceptual quality (SSIM).
4.2.3 Runtime Efficiency Comparison
Table 2 shows a comparison of runtime between NeurOp-Diff and SR3, IDM under different inference steps. Although the runtime of all methods increases with the number of inference steps, our method maintains the fastest execution speed across all tested configurations.
Inference Steps SR3 [saharia2022image] IDM [gao2023implicit] NeurOp-Diff
T=30 1.8 1.47 1.12
T=50 2.16 1.79 1.52
T=100 3.23 2.87 2.61
Table 2: A comparison of the inference time (in seconds) required by different methods to process a single image under varying numbers of inference steps (T = 30, 50, 100).
@@ 4.2.3 Runtime Efficiency Comparison
Table 3: Quantitative comparison of continuous super-resolution (SR) results on UCMerced (PSNR/SSIM/LPIPS).
Method Metric in-distribution out-of-distribution
$2\times$ $3.1\times$ $5.8\times$ $7\times$ $8\times$ $9\times$ $10\times$
SRNO [wei2023super] PSNR 34.32 31.05 27.19 25.73 24.67 24.32 23.92
SSIM 0.786 0.746 0.681 0.649 0.623 0.607 0.592
LPIPS 0.118 0.142 0.194 0.207 0.215 0.221 0.228
IDM [gao2023implicit] PSNR 33.54 30.32 26.37 25.02 23.68 23.34 23.05
SSIM 0.811 0.781 0.692 0.646 0.613 0.601 0.587
LPIPS 0.103 0.119 0.187 0.208 0.219 0.225 0.230
NeurOp-Diff PSNR 33.93 30.76 26.89 25.50 24.51 24.13 23.88
SSIM 0.824 0.792 0.717 0.673 0.650 0.641 0.633
LPIPS 0.098 0.112 0.159 0.188 0.195 0.206 0.212
4.3 Comparison of Continuous SR
Table 4 presents the quantitative comparison of LIIF, SRNO, and NeurOp-Diff on UCMerced. Although LIIF and SRNO achieve higher PSNR, our NeurOp-Diff demonstrates better performance in terms of SSIM. Moreover, in subjective perceptual differences (LPIPS) [zhang2018unreasonable], the images generated by our model show smaller differences compared to the real images. As shown in Figure 4, regression-based continuous super-resolution methods struggle to reconstruct high-frequency details and restore realistic textures, often producing overly smooth images. For instance, fine ground stripes are almost invisible in the images output by SRNO. Even when reconstructing more prominent stripes, the results appear somewhat blurry. In contrast, NeurOp-Diff demonstrates superior capability in restoring these details.
From a quantitative perspective (see Table 3), although SRNO demonstrates slightly better PSNR values at lower magnification factors (e.g., 2x to 3.1x), NeurOp-Diff demonstrates superior performance in terms of SSIM metrics, which better aligns with human visual perception. As the magnification factor increases, the advantages of NeurOp-Diff become more apparent, particularly in challenging scenarios with higher magnification factors (e.g., 7x to 10x), where it consistently maintains higher SSIM values as well as lower LPIPS scores. Visually, NeurOp-Diff exhibits better performance in terms of the clarity of tennis court boundary lines and natural texture transitions in the surrounding environment. These advantages become more pronounced at higher magnification factors. This is primarily because NeurOp-Diff is better at capturing and reconstructing high-frequency information, allowing the generated images at high magnification factors to maintain high visual quality and detail richness. However, SRNO often struggles to fully recover complex texture details when handling larger magnifications, resulting in relatively lacking detail in the generated images.
@@ 4.3 Comparison of Continuous SR
LIIF [chen2021learning] 28.96 0.69 0.182
SRNO [wei2023super] 29.21 0.72 0.164
@@ 4.4 Ablation Studies
To explore the impact of the LR conditional network, we conducted 4× super-resolution ablation experiments on remote sensing datasets. Specifically, we constructed comparison models by replacing the scale-adaptive conditional network in NeurOp-Diff with three different conditional mechanisms: (1) Concatenating the upsampled LR image directly with the ground truth image (SR3 [saharia2022image]); (2) Concatenating LR features encoded by EDSR [Lim2017] with the ground truth image; (3) Concatenating LR features encoded by the neural operator with the ground truth image. As shown in Figure 1, directly using the upsampled LR image as a prior condition results in the least amount of detail recovery. Encoding LR features with EDSR recovers more details, but it still falls short compared to the scale-adaptive network encoded by the neural operator. The quantitative results in Table 5 further validate the effectiveness of the neural operator-based conditioning mechanism, as it achieves the best performance in both PSNR and SSIM. In comparison, it is evident that LR images encoded by the neural operator provide more prior knowledge.
Bicubic EDSR Neural Operator
[…]
Table 5: Comparison of different methods by PSNR and SSIM